# Target Aggregate Data Adjustment Method for Transportability Analysis Utilizing Summary-Level Data from the Target Population

### v1

@@ Abstract
Transportability analysis is a causal inference framework used to evaluate the external validity of randomized clinical trials (RCTs) or observational studies. Most existing transportability analysis methods require individual patient-level data (IPD) for both the source and the target population, narrowing its applicability when only target aggregate-level data (AgD) is available. Besides, accounting for censoring is essential to reduce bias in longitudinal data, yet AgD-based transportability methods in the presence of censoring remain underexplored. Here, we propose a two-stage weighting framework named “Target Aggregate Data Adjustment” (TADA) to address the mentioned challenges simultaneously. TADA is designed as a two-stage weighting scheme to simultaneously adjust for both censoring bias and distributional imbalances of effect modifiers (EM), where the final weights are the product of the inverse probability of censoring weights and participation weights derived using the method of moments. We have conducted an extensive simulation study to evaluate TADA’s performance. Our results indicate that TADA can effectively control the bias resulting from censoring within a non-extreme range suitable for most practical scenarios, and enhance the application and clinical interpretability of transportability analyses in settings with limited data availability.
@@ 1 Introduction
Transportability analysis is a novel causal inference framework used to quantitatively assess the external validity of both randomized and observational studies 1, 2, 3, 4, 5, 6. By evaluating external validity, transportability analyses help bridge the gap between effect estimates observed in a source population and their anticipated impact in the target population 7.
@@
While different transportability methods are available, most require individual participant-level data (IPD) for both the source and target samples 3, 8, 9, 10, 11. Access to such data is often limited, presenting practical challenges to transportability analyses 1. A common scenario is that IPD is only available for the source, while the target population provides aggregate-level data (AgD), typically from disease registries, healthcare databases, or published studies. To address this limitation, recent methodological developments have sought to leverage AgD to improve the feasibility of transportability analyses. For instance, entropy balancing approaches use convex optimization to derive sampling weights that align covariate moments between IPD from the source sample and AgD from the target population, facilitating effect estimate transport. This approach has been refined with calibration techniques, enhancing robustness and accuracy across varied contexts, including randomized clinical trials (RCTs) and observational studies 12, 13. More recently, Quan et al. 14 introduced a double inverse probability weighting (DIPW) method tailored for situations with limited data from the target population, particularly when only regional summary data is available due to privacy constraints. Additionally, some methods incorporate more flexible weighting strategies that use individual-level covariates from the source sample, further improving estimation consistency and efficiency across broader scenarios 15.
@@
While previous studies on these AgD-based transportability methods have explored binary and continuous outcomes, there has been limited exploration for survival analyses critical for oncology. Given the importance of accurately applying trial results to broader populations, recent work has explored different ways to transport survival outcomes from an RCT to target populations. For example, Ramagopalan et al. developed a method for transporting survival estimates: they transported overall survival (OS) effect estimates from US clinical practice to the Canadian population using baseline covariates adjustments based on the Canadian population 16. Zuo et al. applied the proportional hazards and accelerated failure time models to transport survival outcomes in adjuvant colon cancer trials 17. They explored one-trial-at-a-time and leave-one-trial-out methods, demonstrating how varying model structures and moment adjustments can address treatment effect discrepancies due to heterogeneous covariate distributions across trials, providing a structured framework for extending survival results to diverse populations. Lee et al. extended the augmented calibration weighting (ACW) methods with the linear spline-based hazard regression model to generalize treatment effects from the ACTG 175 HIV trial to various target populations without the limitation of proportional hazard (PH) assumption 18.
@@
Notably, censoring in survival analysis critically impacts the performance of traditional methods and presents similar challenges for transportability. Recent efforts have focused on methods to address censoring effects specifically in survival data transportability. One such effort is Berkowitz et al.’s approach, which uses inverse odds of selection weights with survival tools like the Kaplan–Meier estimator and marginal structural Cox models to transport survival outcomes 19. Although effective for estimating counterfactual survival, this method assumes censoring is independent of covariates, limiting its real-world applicability where censoring often depends on covariates. Lee et al. advance this by addressing covariate-dependent censoring in survival transportability 20. Using a doubly robust (DR) estimator, their method integrates inverse probability weighting for sampling, censoring, and treatment assignment with outcome regression, ensuring accuracy even when censoring varies with covariates. Cao et al. expanded on weighting and DR methods to estimate counterfactual survival functions within the target population 21. Unlike Lee et al.’s calibration weighting estimators, this approach employs simpler estimators that rely on direct weight estimation or survival outcome modelling. However, it assumes that the target population data is available through a complex survey with pre-estimated weights, a more idealized scenario than purely aggregate-level data, as it requires well-estimated survey weights for accurate covariate alignment. While these methods have advanced survival outcome transportability, they still largely depend on complete or partial access to individual-level data from the target population, limiting their applicability when only aggregate-level data is available. This limitation highlights a gap in current research, motivating our efforts to enable reliable transportability with covariate-dependent censoring using limited target data.
@@
In this paper, we propose a novel framework called “Target Aggregate Data Adjustment” (TADA) that can account for covariate-dependent censoring when transporting findings to aggregate-level target data. TADA adjusts for potential bias from censoring by incorporating the inverse probability of censoring weights and matches effect modifier distributions between source and target populations using the method of moments (MoM) weighting. TADA then derives final matching weights for each individual in the source dataset through a two-stage scheme, balancing the source and target populations to transport treatment effects.
@@
The paper is organized as follows. In Section 2, we introduce the notations and formalize the causal inference framework for survival outcomes in transportability with necessary assumptions. Section 3 introduces the methodological approaches used in TADA. Sections 4 and 5 cover the simulation settings and present the results, respectively. A discussion is provided in Section 6, with concluding remarks in Section 7. A brief review of existing transportability methods for aggregate target data is provided in the Supplementary Materials.
@@ 2 Notation, Framework and Assumptions
Suppose we are interested in comparing the effectiveness of two treatments. We have access to the source sample $\mathcal{S}$ of size $n_{s}$, with individual data. Let $\textbf{X}\in\mathcal{X}$ be a vector of baseline covariates including effect modifiers for the individuals in the trial and those in the target population. $A\in\{0,1\}$ is the treatment indicator, with $A=1$ referring to active treatment and $A=0$ representing control. $S$ is defined as the binary trial participation indicator where $S=1$ stands for the individual participants in the RCT and $S=0$ otherwise. $T$ is the survival time, defined as the time of event occurrence. Following the potential outcomes framework 22, 23, let $T^{(a)}$ be the potential survival time if a subject receives the treatment $A=a$. Under the stable unit treatment value assumption (SUTVA) 22, the survival time is indicated as $T=T^{(1)}A+T^{(0)}(1-A)$. Let $C$ be the censoring time. In the presence of right censoring, the survival time $T_{i}$ is not always observed for all subjects; instead, we observe $U_{i}=T_{i}\wedge C_{i}$ and $\Delta=I(T\leq C)$, where $\wedge$ represents the minimum of two values, and $I(\cdot)$ is an indicator function. In summary, from the RCT source sample, we observe $\{U_{i},\Delta_{i},A_{i},\textbf{X}_{i},S_{i}=1\}$ from $i=1,\cdots,n_{s}$ subjects.
@@
The target population, $\mathcal{T}$, contains $n_{t}$ individuals. We do not observe either the treatment or outcome but only have access to the sample moments $\bar{h}_{k,\mathcal{T}}=\frac{1}{n_{t}}\sum_{i\in\mathcal{T}}h_{k}(X_{i})$ of certain covariate functions $h_{k}:\mathcal{X}\rightarrow\mathbb{R}$, where $k=1,\dots,K_{h}$.
@@
Let $S_{a}(t)=\Pr(T^{(a)}\geq t)$ be the survival function and $\lambda_{a}(t)=\text{lim}_{h\rightarrow 0}\frac{1}{h}\frac{\Pr(t\leq T^{(a)}\leq t+h)}{\Pr(T^{(a)}\geq t)}$ be the hazard function. Under the proportional hazards assumption, i.e., $\lambda_{1}(t)/\lambda_{0}(t)$ being a constant, the estimand of interest is chosen as the marginal hazard ratio, derived by estimating the causal effect of treatment on survival time within a weighted Cox proportional hazards (PH) model.
@@
In RCTs, treatment assignment $A$ is independent of covariates X given study participation, thus the treatment propensity score $\pi_{A}$ is defined as the fixed probability of receiving the treatment. Let $\pi_{S}(\textbf{X})=\Pr(S=1\mid\textbf{X})$ be the sampling score. We make the following assumptions for the identification of the estimand of interest:
Assumption 1. (Randomization). $\{T^{(0)},T^{(1)}\}\perp\!\!\!\perp A\mid S=1$; and $10<\pi_{A}<$ 1.
@@
Assumption 2. (Ignorability and positivity of trial participation). $\{T^{(0)},T^{(1)}\}\perp\!\!\!\perp S\mid\textbf{X}$; and $10<\pi_{S}(\textbf{X})<$ 1 for all values of X.
@@
Assumption 3. (Noninformative censoring conditional on covariates and treatment). $\{T^{(0)},T^{(1)}\}\perp\!\!\!\perp C\mid(\textbf{X},A,S=1)$, which also implies $T\perp\!\!\!\perp C\mid(\textbf{X},A,S=1)$.
@@
Given the inherent difficulties in directly validating these assumptions, relevant domain knowledge should support their practical plausibility. Assumption 1 holds in the RCT by design. If all information related to trial participation and outcomes is captured in the data, then the first part of Assumption 2 is reasonable. Additionally, the second part of Assumption 2 suggests that missing certain patient characteristics would prevent those patients from participating in the trial 24. Assumption 3 is the common assumption in survival analysis and imposes fewer restrictions compared to the conditional independence assumption between censoring and survival time based solely on treatment 8, 19, 25, 26.
@@ 3.1 Participation Weights: Method of Moments
If IPD is available for both the source population and target population, it is straightforward to obtain participation weights with regression-based methods. However, when IPD is not available for the target population, the limited information provided by AgD leads to additional challenges in estimation and requires a different approach. To address this issue, we propose a MoM-based approach as a feasible alternative for covariate adjustment.
@@
The MoM method’s original application occurs in indirect treatment comparison (ITC), which also requires a similar covariate matching between two RCTs 27. As one of the earlier methods developed for ITC, MoM was initially applied in standard matching-adjusted indirect comparison (MAIC) and has since become a widely adopted approach in ITC. For each covariate of interest, the MoM approach aligns the sample moments calculated from the IPD of the source population with the corresponding moments from the AgD of the target population. Greater availability of AgD, particularly higher-order target sample moments, reduces bias in the population matching process. For instance, the access to additional summary statistics such as the sample standard deviation, median, proportion, or even higher-order moments (e.g., skewness or kurtosis) for covariates results in more accurate estimators compared to relying solely on the sample mean 28, 29. Our approach is conceptually similar to the strategy introduced by Josey et al. 12 where entropy balancing was employed to achieve the same goal. The existing literature highlights the connection between entropy balancing and the MoM approach, particularly in applications to ITC and transportability analyses. Both methods are statistically equivalent, sharing common advantages and robustness 30, while MoM has the advantage of being relatively intuitive in clinical interpretation, discussed in the Supplementary Materials.
@@
Here, we illustrate the theoretical process of participation weights estimation by the MoM approach via first-order sample moments (i.e., sample mean) matching between two populations. Consider covariates X available in both source and target population. Assume that we have the availability of IPD for X in the source sample but just AgD for X in the target sample. For each individual $i$, let $X_{i}\in\textbf{X}_{i}$ be the available IPD in the source sample. Let $\bar{X}$ be the available AgD summary statistic in the target sample.
Assume that the weight for individual $i$ follows the logistic regression:
$w_{i}^{\textbf{MoM}}=\exp(\alpha+X_{i}\cdot k).$
Then $w_{i}^{\textbf{MoM}}$ can be estimated by matching $\bar{X}$ of the target sample with the weighted source sample moment as follows:
$\frac{\sum_{i=1}^{N}X_{i}\cdot w_{i}^{\textbf{MoM}}}{\sum_{i=1}^{N}w_{i}^{\textbf{MoM}}}=\bar{X},$
where the coefficients $\alpha$ and $k$ in the weights will be estimated to satisfy the equality constraint.
Without loss of generality, it can be assumed that $\bar{X}=0$ in theoretical formula (e.g., we could transform/center baseline characteristics in both trials by subtracting $\bar{X}$ in format) and derive the procedure as follows:
$\frac{\sum_{i=1}^{N}X_{i}\cdot w_{i}^{\textbf{MoM}}}{\sum_{i=1}^{N}w_{i}^{\textbf{MoM}}}$ $=\bar{X}$
$\sum_{i=1}^{N}X_{i}\cdot w_{i}^{\textbf{MoM}}$ $=\bar{X}\cdot\sum_{i=1}^{N}w_{i}^{\textbf{MoM}}$
$\sum_{i=1}^{N}X_{i}\cdot w_{i}^{\textbf{MoM}}$ $=0=Q^{\prime}(k),$
where $Q(k)=\sum_{i=1}^{N}w_{i}^{\textbf{MoM}}=\sum_{i=1}^{N}\exp(\alpha+X_{i}\cdot k)$. Since $Q(k)$ is convex, any solution of the equality constraint is unique and corresponds to the global minimum of $Q(k)$. We can follow the process to eventually get the estimates of $w_{i}^{\textbf{MoM}}$.

### v2

@@ Abstract
Transportability analysis is a causal inference framework used to evaluate the external validity of studies by transporting treatment effects from a study sample to an external target population by adjusting for differences in the distributions of their effect modifiers. Most existing methods require individual patient-level data (IPD) for both the source and the target population, narrowing its applicability when only target aggregate-level data (AgD) are available. For survival analysis, accounting for censoring may be needed to reduce bias, yet AgD-based transportability methods in the presence of informative-censoring remain underexplored. Here, we propose a two-stage weighting framework named “Target Aggregate Data Adjustment” (TADA) that can simultaneously adjust for both censoring bias and distributional imbalances of effect modifiers. In our framework the final weights are the product of the time-varying inverse probability of censoring weights and participation weights derived using the method of moments. We have conducted an extensive simulation study to evaluate TADA’s performance. We have applied our methods to a real case study on the squamous non-small-cell lung cancer trial (NCT00981058). Our results indicate that TADA can effectively control the bias resulting from moderate censoring representative of most practical scenarios, and enhance the application and clinical interpretability of transportability analyses in settings with limited data availability.
@@ 1 Introduction
Transportability analysis is a novel causal inference framework used to quantitatively assess the external validity of randomized or observational studies 1, 2, 3, 4, 5, 6. Transportability analysis methods “transport” findings from a source sample, gathered from a randomized clinical trial (RCT) or non-randomized (observational) study to an external target population by adjusting for differing distributions of effect modifiers between the source and target samples 2, 3, 4, 5, 6. By evaluating external validity, transportability analyses help bridge the gap between effect estimates observed in a source population and their anticipated impact in a target population of interest 7.
@@
While different transportability methods are available, most require individual participant-level data (IPD) for both the source and target samples 3, 8, 9, 10, 11. Access to such data is often limited, presenting practical challenges to transportability analyses 1. A common scenario is that IPD is only available for the source study, while the target population data is limited to aggregate-level data (AgD) available from publications. Recent methods developments have sought to leverage AgD to improve the feasibility of transportability analyses. For instance, entropy balancing approaches use convex optimization to derive sampling weights that align covariate moments between IPD from the source sample and AgD from the target population, facilitating effect estimate transport. This approach has been refined with calibration techniques, including an exponential-tilting calibration scheme which achieves semiparametric efficiency under mild regularity conditions 12, 13. In another study by Quan et al. 14, they introduced a double inverse probability weighting (DIPW) method tailored for situations where only the summary data is available due to privacy constraints of specific countries/regions. Other proposed methods incorporate more flexible weighting strategies that use individual-level covariates from the source sample to seeks additional covariate balance between the treated and control groups 15.
@@
Existing studies on these AgD-based transportability methods have explored binary and continuous outcomes, and there has been limited exploration for time-to-event outcomes. Given the importance of accurately applying trial results to broader populations, recent work has explored different ways to transport survival outcomes from an RCT to target populations. For example, Ramagopalan et al. 16 developed a method for transporting survival estimates: they transported overall survival (OS) effect estimates from US clinical practice to the Canadian population using baseline covariates adjustments based on the Canadian population, but did so with individual-level data available in both populations. Zuo et al. 17 applied proportional hazards and accelerated failure time models to transport survival outcomes in adjuvant colon cancer trials. They explored one-trial-at-a-time and leave-one-trial-out methods, demonstrating how varying model structures and moment adjustments can address treatment effect discrepancies due to heterogeneous covariate distributions across trials, providing a structured framework for extending survival results to diverse populations. However, the framework likewise presumes the availability of individual-level data for both the source and each target trial. Lee et al. 18 extended the augmented calibration weighting (ACW) methods with the linear spline-based hazard regression model to generalize treatment effects from the ACTG 175 HIV trial to various target populations without the limitation of proportional hazard (PH) assumption. This extension requires target-population IPD (or an equivalent synthetic micro-dataset), imposing the same practical constraint as mentioned.
@@
Informative censoring in survival analysis can critically impact the performance of statistical methods. For transportability analyses, this is no different. There has been recent work developing methods to account for censoring when performing transportability analysis on survival data. One such work is by Berkowitz et al. 19, which uses inverse odds of selection weights with survival tools like the Kaplan–Meier estimator and marginal structural Cox models to transport survival outcomes. Although effective for estimating counterfactual survival times, this method assumes censoring is independent of covariates, potentially limiting its real-world applicability. Lee et al. 20 advance this by addressing covariate-dependent censoring in survival transportability. Using a doubly robust (DR) estimator, their method integrates inverse probability weighting for sampling, censoring, and treatment assignment with outcome regression, ensuring accuracy even when censoring depends on covariates. Cao et al. 21 expanded on weighting and DR methods to estimate counterfactual survival functions within the target population. Unlike Lee et al.’s calibration weighting estimators, this approach employs simpler estimators that rely on direct weight estimation or survival outcome modelling. These methods have advanced survival outcome transportability, but they still require access to IPD of the target population. There is a current gap in transportability analysis methods that can account for covariate-dependent censoring in settings where only AgD is available in the target population.
@@
In this paper, we propose a novel framework called “Target Aggregate Data Adjustment” (TADA) that can account for covariate-dependent time-varying censoring when transporting findings to target AgD. Our TADA framework can incorporate inverse probability of censoring weights with the method of moments (MoM) weighting to account for differences in effect modifier distributions between source and target populations. TADA then derives final matching weights for each individual in the source dataset through a two-stage scheme, balancing the source and target populations to transport treatment effects. The R codes to implement TADA and reproduce the simulation results can be obtained from https://github.com/CoreClinicalSciences/TADA. TADA is also available as a main module in TransportHealth, a recently published open-source R package developed by Park et al. 22 for transportability analyses and causal reasoning.
@@
The paper is organized as follows. In Section 2, we introduce the notations and formalize the causal inference framework for survival outcomes in transportability with necessary assumptions. Section 3 introduces the methodological approaches used in TADA. Sections 4 and 5 cover the simulation settings and present the results, respectively. Sensitivity analyses are provided in Section 6. In Section 7, we apply TADA to transport the Kaplan–Meier survival curve of a control arm from the stage IV squamous non-small-cell lung cancer (NSCLC) treatment with the inhibitor of EGF receptor (SQUIRE) trial (NCT00981058) 23 to a real-world target population with stage IV NSCLC in Ontario, Canada 24. A discussion is provided in Section 8, with concluding remarks in Section 9. In the Supplementary Materials, we provide a brief review of existing transportability methods for aggregate target data, a theoretical justification of unbiased estimation of TADA and additional simulation results.
@@ 2 Notation, Framework and Assumptions
Suppose we are interested in comparing the effectiveness of two treatments. We have access to the source sample $\mathcal{S}$ of size $N_{\textit{s}}$, with individual data. Let $\textbf{X}\in\mathcal{X}$ be a vector of baseline covariates, consisting of two subsets: (1) covariates $\mathbf{L}$, which include variables associated with the risk of being censored (e.g., dropout or loss to follow-up) and required to ensure conditional exchangeability of counterfactual survival times and censoring; and (2) effect modifiers $\mathbf{E}$, which include variables that influence how treatment effects may vary across different populations and required to ensure conditional exchangeability of counterfactual survival times and trial participation. For simplicity and clarity in presentation, we uniformly denote both subsets together as $\mathbf{X}$ throughout the manuscript. Let $A\in\{0,1\}$ be the treatment indicator, with $A=1$ referring to active treatment and $A=0$ representing control. Let $S\in\{0,1\}$ be the trial participation indicator, with $S=1$ referring to trial participation in the RCT, and $S=0$ otherwise. Let $T$ denote the survival time, defined as the event occurrence time. Following the potential outcomes framework 25, 26, let $T^{(a)}$ be the potential survival time if a subject receives the treatment $A=a$. Under the stable unit treatment value assumption (SUTVA) 25, the survival time is indicated as $T=T^{(1)}A+T^{(0)}(1-A)$. Let $C$ be the censoring time. In the presence of right censoring, the survival time $T_{i}$ is not always observed for all subjects; hence, we define observed time as $U_{i}=T_{i}\wedge C_{i}$ and censoring indicator $\Delta_{i}=I(T_{i}\leq C_{i})$, where $\wedge$ represents the minimum of two values, and $I(\cdot)$ is an indicator function. In summary, from the RCT source sample, we observe $\{U_{i},\Delta_{i},A_{i},\textbf{X}_{i},S_{i}=1\}$ from $i=1,\cdots,N_{\textit{s}}$ subjects.
@@
The target population, $\mathcal{T}$, contains $N_{\textit{t}}$ individuals. We do not observe either the treatment or outcome but have access to the sample moments $\overline{h}_{k,\mathcal{T}}=\frac{1}{N_{\textit{t}}}\sum_{i\in\mathcal{T}}h_{k}(\mathbf{X}_{i})$ of certain covariate functions $h_{k}:\mathcal{X}\rightarrow\mathbb{R}$, where $k=1,\dots,K$.
@@
Let $S_{a}(t)=\Pr(T^{(a)}\geq t)$ be the survival function and $\lambda_{a}(t)=\text{lim}_{h\rightarrow 0}\frac{1}{h}\frac{\Pr(t\leq T^{(a)}\leq t+h)}{\Pr(T^{(a)}\geq t)}$ be the hazard function. Under the proportional hazards assumption, i.e., $\lambda_{1}(t)/\lambda_{0}(t)$ being a constant, the estimand of interest is chosen as the marginal hazard ratio, derived by estimating the causal effect of treatment on survival time within a weighted Cox PH model.
@@
In RCTs, treatment assignment $A$ is independent of covariates X given study participation $S$, thus the treatment propensity score $\pi_{A}$ is defined as the fixed probability of receiving the treatment. Let $\pi_{S}(\textbf{X})=\Pr(S=1\mid\textbf{X})$ be the sampling score. We make the following assumptions for the identification of the estimand of interest:
Assumption 1. (Randomization) $\{T^{(0)},T^{(1)}\}\perp\!\!\!\perp A\mid S=1$; and $10<\pi_{A}<$ 1.
@@
Assumption 2. (Ignorability and positivity of trial participation) $\{T^{(0)},T^{(1)}\}\perp\!\!\!\perp S\mid\textbf{X}$; and $10<\pi_{S}(\textbf{X})<$ 1 for all values of X.
@@
Assumption 3. (Noninformative censoring conditional on covariates and treatment) $\{T^{(0)},T^{(1)}\}\perp\!\!\!\perp C\mid(\textbf{X},A,S=1)$, which also implies $T\perp\!\!\!\perp C\mid(\textbf{X},A,S=1)$.
@@
Given the inherent difficulties in directly validating these assumptions, relevant domain knowledge should support their practical plausibility. Assumption 1 holds in the RCT by design. If all information related to trial participation and outcomes is captured in the data, then the first part of Assumption 2 is reasonable. Additionally, the second part of Assumption 2 requires that all subgroups defined by covariates have a non-zero probability of trial participation. Specifically, if certain patient characteristics are completely unobserved or excluded by design, then individuals with these characteristics would have no possibility of participating in the trial, violating the positivity condition and limiting the transportability of trial results to such populations 27. Assumption 3 is a common assumption in survival analysis and imposes fewer restrictions compared to the conditional independence assumption between censoring and survival time based solely on treatment 8, 19, 28, 29.
@@ 3.1 Participation Weights: Method of Moments
When only target AgD is available, conducting transportability analysis faces challenges due to limited information that is provided by AgD. To address this issue, we propose a MoM-based approach as a feasible alternative for covariate adjustment.
@@
The original application of MoM methods was in matching-adjusted indirect comparison (MAIC) to achieve balance of covariates between two RCTs of comparison 30. For each covariate of interest, the MoM approach aligns the sample moments calculated from the IPD of the source population with the corresponding moments from the AgD of the target population. Greater availability of AgD, particularly higher-order target sample moments, reduces bias in the population matching process. For instance, the access to additional summary statistics such as the sample standard deviation, median, proportion, or even higher-order moments (e.g., skewness or kurtosis) for covariates can result in more accurate estimators compared to relying solely on the sample mean 15. Our approach is conceptually similar to the strategy introduced by Josey et al. 12 where entropy balancing was employed to achieve the same goal. The existing literature highlights the connection between entropy balancing and the MoM approach, particularly in applications to ITC and transportability analyses. Both methods are statistically equivalent, sharing common advantages and robustness 31, while MoM has the advantage of being relatively intuitive in clinical interpretation, as discussed in section S1 of Supplementary Materials.
@@
Here, we illustrate the theoretical process of participation weight estimation by the MoM approach through matching first-order sample moments (sample means). For simplicity and ease of illustration, we describe the process for a single covariate component, while the approach naturally extends to multiple covariates. Consider a specific covariate component $X$ available in both the source and target populations. Assume that we have IPD for $X$ in the source sample, but only the sample mean $\overline{X}$ for this covariate in the target sample.
For each individual $i$, let $X_{i}$ represent the individual-level covariate data available in the source sample. We construct participation weights in the widely used exponential tilting form 12, 15, 32, 33. Formally, we model participation weights for individual $i$ as:
$w_{i}^{\text{MoM}}=\exp\bigl{(}\gamma+X_{i}\cdot\zeta\bigr{)},$
where $\gamma$ is a normalizing constant and $\zeta$ is a parameter to be determined such that the weighted source sample mean matches the target sample mean. Then MoM-based method matches the weighted source sample mean with the target sample mean:
$\frac{\sum_{i=1}^{N_{\textit{s}}}X_{i}\cdot w_{i}^{\text{MoM}}}{\sum_{i=1}^{N_{\textit{s}}}w_{i}^{\text{MoM}}}=\overline{X}.$
Without loss of generality, we apply a transformation by centering the covariate $X$ in the source population, using the target population mean $\overline{X}$. This yields a new covariate $X_{i,\text{CENTERED}}=X_{i}-\overline{X}$, resulting in the simplified condition where $\overline{X}_{\text{CENTERED}}=0$ in the target population and naturally cancels out the intercept term $\gamma$ without separate estimation. Thus, the weights simplify to:
$w_{i}^{\text{MoM}}=\exp\bigl{(}X_{i,\text{CENTERED}}\cdot\zeta\bigr{)},$
and the equality constraint becomes:
$\frac{\sum_{i=1}^{N_{\textit{s}}}X_{i,\text{CENTERED}}\cdot w_{i}^{\text{MoM}}}{\sum_{i=1}^{N_{\textit{s}}}w_{i}^{\text{MoM}}}=0.$
By further simplifying the constraint as $\sum_{i=1}^{N_{\textit{s}}}X_{i,\text{CENTERED}}\cdot w_{i}^{\text{MoM}}=0=Q^{\prime}(\zeta)$, where the objective function $Q(\zeta)$ is given by $Q(\zeta)=\sum_{i=1}^{N_{\textit{s}}}\exp(X_{i,\text{CENTERED}}\cdot\zeta)$, the solution for $\zeta$ that satisfies this equality constraint is unique and globally optimal due to the convexity of $Q(\zeta)$. Thus, stable estimates of participation weights are obtained.